### v2

# Lower semicontinuity, Stoilow factorization and principal maps

Abstract.
We consider a strengthening of the usual quasiconvexity condition of Morrey in two dimensions, which allows us to prove lower semicontinuity for functionals which are unbounded as the determinant vanishes. This notion, that we call principal quasiconvexity, arose from the planar theory of quasiconformal mappings and mappings of finite distortion. We compare it with other quasiconvexity conditions that have appeared in the literature and provide a number of concrete examples of principally quasiconvex functionals that are not polyconvex. The Stoilow factorization, that in the context of maps of integrable distortion was developed by Iwaniec and Šverák, plays a prominent role in our approach.
Dedicated to Vladimir Šverák on the occasion of his 65th birthday
1. Introduction
Given a real $n\times n$ matrix $A$, its (outer) distortion is defined by
$K_{A}\equiv\left\{\begin{array}[]{cl}|A|^{n}/\det A&\mbox{ if }\det A>0,\\ 1&\mbox{ if }A=0,\\ +\infty&\mbox{ if }A\neq 0\mbox{ and }\det A\leq 0,\end{array}\right.$
where $|A|\equiv\max_{x\in\mathbb{S}^{n-1}}|Ax|$ and $|Ax|$ is the usual euclidean norm of $Ax\in\mathbb{R}^{n}$. When $\Omega$ is an open subset of $\mathbb{R}^{n}$ and $u\colon\Omega\to\mathbb{R}^{n}$ is locally of Sobolev class $\mathrm{W}^{1,1}$ in $\Omega$, denoted $u\in\mathrm{W}^{1,1}_{\operatorname{loc}}(\Omega,\mathbb{R}^{n})$, the distortion function $K_{u}(x)\equiv K_{\textup{D}u(x)}$ is an $\mathscr{L}^{n}$-almost everywhere defined extended real-valued function on $\Omega$. If the distortion $K_{u}<+\infty$ almost everywhere and the Jacobian $J_{u}\equiv\det\textup{D}u\in\mathrm{L}^{1}_{\operatorname{loc}}(\Omega)$, then $u$ is called a mapping of finite distortion, and if furthermore $K_{u}\in\mathrm{L}^{\infty}(\Omega)$, then $u$ is a weakly quasiregular map. A striking result of Reshetnyak states that a quasiregular map, so a map $u\in\mathrm{W}^{1,n}_{\operatorname{loc}}(\Omega,\mathbb{R}^{n})$ with $K_{u}\in\mathrm{L}^{\infty}(\Omega)$, admits a representative that is continuous, open and discrete. In the two-dimensional case, $n=2$, which is our focus here, this result can be further refined via the so-called Stoilow factorization, which asserts that for an open subset $\Omega$ of $\mathbb{C}$ a map $f\colon\Omega\to\mathbb{C}$ is continuous, open and discrete if and only if $f=h\circ F$, where $F\colon\mathbb{C}\to\mathbb{C}$ is a homeomorphism and $h\colon F(\Omega)\to\mathbb{C}$ is holomorphic (see [43], and [35] for a modern exposition).
In [31] Tadeusz Iwaniec and Vladimir Šverák brought together ideas from Geometric Function Theory and Nonlinear Elasticity proving, again in two dimensions, that if $u\in\mathrm{W}^{1,2}_{\operatorname{loc}}(\Omega)\equiv\mathrm{W}^{1,2}_{\operatorname{loc}}(\Omega,\mathbb{C})$ has distortion function $K_{u}\in\mathrm{L}^{1}(\Omega)$, then (the precise representative of) $u$ is continuous, open and discrete (see also [5] for various refinements). The aim of this note is to explain the role of Stoilow factorization, and in particular of the result of Iwaniec and Šverák, in an apparently unrelated context, namely in proving lower semicontinuity results in situations where the energy functionals do not satisfy the standard growth assumptions, such as in typical variational models of hyperelasticity or variational problems arising in Geometric Function Theory.
We briefly review some background results from the calculus of variations before elaborating further on the geometric function theory concepts, such as Stoilow factorization, the distortion function and the principal maps.
First recall that when $\Omega$ is a bounded open subset of $\mathbb{R}^{n}$ and the exponent $p\geq 1$, then the $\mathrm{W}^{1,p}$-quasiconvexity of a functional $\mathbf{E}\colon\mathbb{R}^{n\times n}\to\mathbb{R}$, meaning that
${\bf E}(A)\leq\fint_{\Omega}{\bf E}(\textup{D}u(z))\,\mathrm{d}m(z),\quad\;{\rm for\,all}\;u\in A+\mathrm{W}^{1,p}_{0}(\Omega,\mathbb{R}^{n})\;{\rm and}\;A\in\mathbb{R}^{2\times 2},$
is equivalent to the sequential weak lower semicontinuity on $\mathrm{W}^{1,p}(\Omega,\mathbb{R}^{n})$ (briefly $\mathrm{W}^{1,p}$-swlsc) of the corresponding variational integral
(1.1) $\mathscr{E}[u]\equiv\int_{\Omega}\!{\bf E}(\textup{D}u)\,\mathrm{d}x,$
provided that ${\bf E}$ is non-negative and in addition has standard $p$-growth:
(1.2) $|\mathbf{E}(A)|\leq c\bigl{(}|A|^{p}+1\bigr{)}\quad\forall\,A\in\mathbb{R}^{n\times n}.$
It is easy to see that this equivalence persists if the functional $\mathbf{E}$ satisfies (1.2) but, instead of being non-negative, satisfies instead the condition
(1.3) $\liminf_{|A|\to+\infty}\frac{\mathbf{E}(A)}{|A|^{p}}\geq 0.$
However, when $\mathbf{E}$ has only the standard $p$-growth (1.2) for an exponent $p>1$, then the variational integral (1.1) can fail to be $\mathrm{W}^{1,p}$-swlsc, due to possible concentration effects at the boundary [38, 14, 36, 1] (this happens already for $\mathbf{E}(A)=\det A$, $A\in\mathbb{R}^{2\times 2}$, and for $n=p=2$, see [23]).
On the other hand, as observed by Meyers [37], the result persists if the variational integral $\mathscr{E}$ is restricted to Dirichlet classes: if $g\in\mathrm{W}^{1,p}(\mathbb{R}^{n},\mathbb{R}^{n})$, then for a functional $\mathbf{E}$ satisfying (1.2), the variational integral $\mathscr{E}$ is swlsc on $g+\mathrm{W}^{1,p}_{0}(\Omega,\mathbb{R}^{n})$ if and only if $\mathbf{E}$ is $\mathrm{W}^{1,p}$-quasiconvex. We refer to [17] for further discussions and to [21] for some related results also linking quasiconvexity to coercivity of $\mathscr{E}$ in Dirichlet classes.
In the absence of the standard $p$-growth condition (1.2), even for the Dirichlet classes $g+\mathrm{W}^{1,p}_{0}(\Omega,\mathbb{R}^{n})$ the sufficiency of $\mathrm{W}^{1,p}$-quasiconvexity for swlsc of $\mathscr{E}$ becomes unclear (in fact, counterexamples exist if the functional $\mathbf{E}$ is allowed to be discontinuous and extended real-valued), but the necessity of $\mathrm{W}^{1,p}$-quasiconvexity for $\mathrm{W}^{1,p}$-swlsc remains true [14].
In this situation one introduces a strengthening of the $\mathrm{W}^{1,p}$-quasiconvexity condition, the so-called closed $\mathrm{W}^{1,p}$-quasiconvexity. This reinstates the sufficiency for $\mathrm{W}^{1,p}$-swlsc of $\mathscr{E}$ on $\mathrm{W}^{1,p}(\Omega,\mathbb{R}^{n})$, provided the functional $\mathbf{E}$ satisfies also (1.3). But then the necessity becomes unclear (in fact, when $\mathbf{E}$ is allowed to be lower semicontinuous and extended real-valued, the necessity of closed $\mathrm{W}^{1,p}$-quasiconvexity fails, see [34]).
These issues become particularly relevant in dealing with variational problems where the effective domain of the energy functional is the set
(1.4) +$\mathbb{R}^{2\times 2}_{+}\equiv\bigl{(}\mathbb{R}^{2\times 2}\cap\{\det>0\}\bigr{)}\cup\{0\},$
or in other words, the set of matrices $A$ with finite distortion $K_{A}$. Such functionals arise naturally for instance in elasticity, where e.g. the classical second invariant, the sum of the singular values divided by $\det(A)$, is given by
(1.5) $I_{2}(A)=K_{A}+\frac{1}{K_{A}},\qquad A\in\mathbb{R}^{2\times 2}_{+}.$
However, two central aspects of the theory in this context remain somewhat obscure. Firstly, it is not clear how ${\bf E}(A)$ should diverge when $\det(A)$ tends to 00, to have a feasible theory for the corresponding variational integral $\mathscr{E}$. Secondly, it seems desirable to rule out pathological deformations (for example homeomorphisms whose determinant is zero a.e or sequences of maps which converge to such maps [28, 24]). Thus it is not obvious what are the appropriate function spaces in which minimisers should belong to, while experiments do not help to clarify these issues as we are close to a regime where the elasticity theory breaks down. For these reasons, and since the general structure of quasiconvex functionals is far from clear, the functionals defined on +$\mathbb{R}^{n\times n}_{+}$ that have been succesfully used in the elasticity literature, have mostly been polyconvex. See for example, [13, 22, 26] and particularly [11, 12] for the setting and success of polyconvexity.
Similar obstacles were faced in our recent study [4] of the local Burkholder functional $\mathbf{B}_{K}$, see (1.8) and Subsection 1.2 for the details. This functional lives on the $K$-quasiconformal well
(1.6) $Q_{2}(K)\equiv\{A\in\mathbb{R}^{2\times 2}:\,K_{A}\leq K\},$
while outside the well it is +$\infty$+ ∞, $\mathbf{B}_{K}(A)=+\infty$ when $A\notin Q_{2}(K)$. Because of sign of $\mathbf{B}_{K}$ on $Q_{2}(K)$ it is easy to see that the closed $\mathrm{W}^{1,p}$-quasiconvexity of $\mathbf{B}_{K}$ is a necessary condition for $\mathrm{W}^{1,p}$-swlsc of the associated variational integral. On the other hand, the sufficiency is now non-obvious. This is the opposite situation of what happens for non-negative extended real-valued functionals.
Such viewpoints apply to all dimensions, but for a planar homeomorphism an application of the chain rule yields the particular relation
$\int_{\Omega}\!K_{f}(z)\,\mathrm{d}m(z)=\int_{f(\Omega)}\!|\textup{D}f^{-1}(w)|^{2}\,\mathrm{d}m(w),$
bridging mappings of integrable distortion with the bi-Sobolev maps. Recalling (1.5), this gives yet another argument in favour of studying the $\mathrm{W}^{1,2}$-mappings with $K\in\mathrm{L}^{1}(\Omega)$.
1.1. Principal maps
To deal with the issues on $\mathrm{W}^{1,p}$-swlsc described above, we next turn to the role of Stoilow factorization in the setting of lower semicontinuity. Here the key notion is that of a principal map:
Definition 1.1.
A principal map is an orientation preserving homeomorphism $f\colon\mathbb{C}\to\mathbb{C}$ of class $\mathrm{W}^{1,1}_{\operatorname{loc}}(\mathbb{C})$, with $K_{f}\in\mathrm{L}^{1}(\mathbb{D})$, which is conformal outside the unit disc $\mathbb{D}$ and has the Laurent expansion
$f(z)=b_{0}z+\frac{b_{1}}{z}+\phi(z),\quad{\rm where}\quad|b_{1}|<|b_{0}|\quad{\rm and}\quad\phi(z)\equiv\sum_{j=2}^{\infty}\frac{b_{j}}{z^{j}}\;\mbox{ for }\,|z|>1.$
Since $1/z=\bar{z}$ for $|z|=1$, we have $f(z)=A_{f}(z)+\phi(z)$ for $|z|=1$, where the linear map $A_{f}\in\mathbb{R}^{2\times 2}$ (that we identify with its matrix) has in the complex notation the form $A_{f}(z)\equiv b_{0}z+b_{1}\bar{z}$. In particular, $A_{f}$ has positive determinant. We also record that
$\fint_{\mathbb{D}}\!\textup{D}f(z)\,\mathrm{d}m(z)=A_{f}.$
Notice that it is standard to normalize principal maps to have $b_{0}=1$. We shall refer to such maps as normalized principal maps. It is useful to recall that the class of all normalized $\mathrm{W}^{1,2}$-principal maps forms a normal family (see [5]); in fact, any sequence of such normalized principal maps admits a subsequence that converges uniformly on $\mathbb{C}$ to a normalized principal map. On the other hand, principal maps with the mere $\mathrm{W}^{1,1}_{\operatorname{loc}}$-regularity do not form a normal family; a sequence of such Sobolev homeomorphisms can have a BV-limit.
We also recall that the inverse of a principal map lies in $\mathrm{W}^{1,2}_{\operatorname{loc}}(\mathbb{C})$ and it has finite distortion, c.f. [28].
Definition 1.2.
A functional ${\bf E}\colon\mathbb{R}^{2\times 2}_{+}\to\mathbb{R}\cup\{+\infty\}$ is called $\mathrm{W}^{1,p}$-principal quasiconvex if it satisfies the Jensen inequality for all principal maps of class $\mathrm{W}^{1,p}_{\operatorname{loc}}(\mathbb{C})$. More precisely, given a principal map $f\colon\mathbb{C}\to\mathbb{C}$ of class $\mathrm{W}^{1,p}_{\operatorname{loc}}(\mathbb{C})$, we require that
(1.7) ${\bf E}(A_{f})\leq\fint_{\mathbb{D}}\!{\bf E}(\textup{D}f(z))\,\mathrm{d}m(z).$
Notice that if $f\in\mathrm{C}^{1}(\mathbb{D})\cap\mathrm{W}^{1,1}(\mathbb{D})$ with $\det\textup{D}f>0$, and for some linear $A\in\mathbb{R}^{2\times 2}_{+}$ we have $f-A\in\mathrm{W}^{1,1}_{0}(\mathbb{D})$, so $f$ is a classical test function for the quasiconvexity inequality, then $f$ can be extended as principal map off the unit disc. For example if $A(z)=b_{0}z+b_{1}\bar{z}$, then the map $f^{\rm prin}(z)=f(z)\mathbf{1}_{\mathbb{D}}+\mathbf{1}_{\mathbb{C}\setminus\mathbb{D}}\bigl{(}b_{0}z+\frac{b_{1}}{z}\bigr{)}$ is principal in the sense of Definition 1.1. Therefore principal quasiconvexity implies a standard quasiconvexity inequality for sufficiently smooth maps. However, from the classical area formula we that $-\det(A)$ is principal quasiconvex, while the determinant itself is not. That is, quasiconvexity does not imply principal quasiconvexity in general.
On the other hand, the class of boundary values of all principal maps is much larger than just the linear ones. In fact, up to a Möbius transform, the class consists of the traces on $\partial\mathbb{D}$ of conformal maps of $\mathbb{C}\setminus\overline{\mathbb{D}}$ onto Jordan domains $\Omega$, with $\infty\in\Omega$ and boundary $\partial\Omega$ sufficiently regular.
Let us then briefly explain why principal quasiconvexity is relevant in dealing with variational integrals. The proof of Morrey [38] for the lower semicontinuity of functionals with standard quasiconvex integrands goes basically as follows. Firstly, given a weakly converging sequence $\{\textup{D}\psi_{j}\}$ of Sobolev functions, one localises the sequence to achieve convergence to a linear map. Second, one replaces this localized sequence with one of affine boundary values, $u_{j}-Az\in\mathrm{W}_{0}^{1,2}(\Omega)$, via a carefully chosen cutoff. And third, one applies directly the definition of quasiconvexity to each such $u_{j}$.
However, such a strategy breaks down for functionals defined on +$\mathbb{R}^{2\times 2}_{+}$ because the cut-off modification might give negative determinants on sets of positive measure.
We hence approach the problem via other methods, namely gradient Young measures with principal quasiconvexity. In fact, the gradient Young measures which the weakly convergent sequences generate often provide a convenient and flexible context for localisation arguments; we refer for the definitions and basic properties of Young measures to the monographs [39, 41, 42]. In particular, we will use the notation and terminology of gradient Young measures as described in our paper [4].
The use of gradient Young measures also enables then new methods and views to analyse the localisation limits. In particular, it is here that the principal quasiconvexity shows its usefulness. For instance, in [3] one first observes that weakly converging sequences of quasiregular mappings generate gradient Young measures that are supported on the quasiconformal wells $Q_{2}(K)$ defined at (1.6). And second, thanks to Stoilow factorization, the properties of those Young measures can be analysed with help of principal maps, via the next
Theorem 1.3.
Consider a gradient Young measure $\nu\in\mathscr{M}^{s}_{\textup{qc}}(Q_{2}(K))$, where $\frac{2K}{K+1}<s<\frac{2K}{K-1}$. Then there is a sequence of $K$-quasiconformal principal maps which generates in $\mathbb{D}$ the Young measure $\nu$.
For mappings of bounded distortion as above, the classical Stoilow factorization is among the main tools in the above result. However, looking for similar methods for mappings of finite or integrable distortion, it is the Iwaniec–Šverák Stoilow factorization that provides us the following important improvement, see [4].
Theorem 1.4.
Let $\nu\in\mathscr{M}^{2}_{\mathrm{qc}}(\mathbb{R}^{2\times 2}_{+})$ be generated by $(\textup{D}\psi_{j})$, where $(\psi_{j})$ is a sequence of homeomorphisms that is bounded in $\mathrm{W}^{1,2}(\mathbb{D})$ and such that $\int_{\mathbb{D}}\!K^{q}_{\psi_{j}}\,\mathrm{d}m(z)\leq C$ for some $q>1$. Then there is a sequence of maps $f_{j}\colon\mathbb{C}\to\mathbb{C}$ such that:
  1. (1)
    $f_{j}$ are principal maps;
  2. (2)
    $(f_{j})\subset\mathrm{W}^{1,2}_{\operatorname{loc}}(\mathbb{D})$ is bounded and $(\textup{D}f_{j}|_{\mathbb{D}(0,R)})$ generates $\nu$ for all $R<1$;
  3. (3)
    $\psi_{j}=h_{j}\circ f_{j}$ for some conformal maps $h_{j}\colon f_{j}(\mathbb{D})\to\psi_{j}(\mathbb{D})$.
In order to apply these ideas we need that ${\bf E}(\textup{D}f_{j})$ is equi-integrable. Thus the proofs of two of our theorems split naturally into two propositions. The first states that the growth conditions, such as given in (1.11) below, guarantee that the sequence ${\bf E}(\textup{D}f_{j})$ is equi-integrable which rules out concentration effects. The second proposition, needed in Theorem 1.6, yields lower semicontinuity once equi-integrability is available, taking care of oscillation effects. The result is satisfactory but equi-integrability is a consequence of our assumption that the distortions are bounded in $\mathrm{L}^{1+\epsilon}$ for some $\epsilon>0$. As a matter of fact in [4] a similar issue arises from the Burkholder functional itself. Fortunately the scope of principal quasiconvexity and Stoilow factorization goes beyond oscillation effects and our third theorem shows that in the case of the Burkholder functional, principal quasiconvexity allows us to use a blow-up technique to prove lower semicontinuity in the borderline case, where equi-integrability cannot be assumed.
We next describe the results.
1.2. Main results
To illustrate the uses of principal quasiconvexity we next present three lower semicontinuity theorems, and start with the classical Burkholder functional $\mathbf{B}_{p}\colon\mathbb{R}^{2\times 2}\to\mathbb{R},$ which for exponents $p\geq 2$ is given by
(1.8) $\mathbf{B}_{p}(A)\equiv\Bigl{(}(\tfrac{p}{2}-1)|A|^{2}-\tfrac{p}{2}\det A\Bigr{)}|A|^{p-2}.$
We note that $\mathbf{B}_{2}=-\det$ while for $p>2$, in terms of the $K$-quasiconformal well (1.6), we have $\mathbf{B}_{p}(A)\leq 0$ if and only if $A\in Q_{2}\left(\frac{p}{p-2}\right)$. Thus, fixing $K\equiv\frac{p}{p-2}$, or $p=\frac{2K}{K-1}$, one is led [4] to consider the local Burkholder functional
$\mathbf{B}_{K}(A)\equiv\left\{\begin{array}[]{ll}\mathbf{B}_{p}(A)&\mbox{ if }A\in Q_{2}(K)\\ +\infty&\mbox{ if }A\in\mathbb{R}^{2\times 2}\setminus Q_{2}(K).\end{array}\right.$
We emphasize the rather unusual feature of $\mathbf{B}_{K}$ that it is nonpositive and $p$-homogeneous on its effective domain $Q_{2}(K)$. As is well known, a planar $K$-quasiregular map lies in $\mathrm{W}^{1,p}_{\operatorname{loc}}$ for every $p<p_{K}\equiv\frac{2K}{K-1}$, while in general the integrability of the derivative fails at the borderline $p_{K}=\frac{2K}{K-1}$. Instead its derivative is locally in the Marcinkiewicz space $\mathrm{L}^{p_{K},\infty}$ (see [2]). On the other hand, a deep result of [6] states that for $K$-quasiregular maps $f\colon\Omega\to\mathbb{C}$, the integral
$0\geq\int_{C}\!{\bf B}_{K}(\textup{D}f)\,\mathrm{d}m(z)>-\infty$
for compact subsets $C$ of $\Omega$. However, even when $K$-quasiregular maps satisfy $f_{j}\rightharpoonup f$ weakly in $\mathrm{W}^{1,p_{K}}$ it still is far from clear whether $\bigl{(}\mathbf{B}_{p_{K}}(\textup{D}f_{j})\bigr{)}$ is locally equi-integrable. Thus the closed quasiconvexity of $\mathbf{B}_{K}$, proven in [4], led there to results that correspond to the $\mathrm{W}^{1,q}$-swlsc of $\mathbf{B}_{K}$ only for $q>p_{K}$. We refer to [19, 20] for the origins of the functional, to [6, 7, 9, 27, 44, 30] for its relation to quasiconvexity and to [5, 15, 16, 18, 40] for results in relation with Iwaniec’s conjecture.
In this note we in particular extend, using principal quasiconvexity and Stoilow factorization, the $\mathrm{W}^{1,q}$-swlsc of the local Burkholder functional up to the borderline case $q=p_{K}$. This is a consequence of a stronger and in a sense more natural result. To explain why the result is natural we fix a bounded open subset $\Omega$ of $\mathbb{C}$ and a $K$-quasiregular map $g\colon\mathbb{C}\to\mathbb{C}$, where $K>1$. Let $p=p_{K}\equiv 2K/(K-1)$ and consider the variational problem of minimizing the Burkholder energy
$\mathcal{B}[v]\equiv\int_{\Omega}\!\mathbf{B}_{p}(\textup{D}v)\,\mathrm{d}m(z)$
over all $K$-quasiregular maps $v\in g+\mathrm{W}^{1,2}_{0}(\Omega)$. First one notes that from [6] (see also below) it follows that the Burkholder energy $\mathcal{B}[v]$ is bounded below over $v\in g+\mathrm{W}^{1,2}_{0}(\Omega)$. Next, let $(u_{j})$ be a minimizing sequence, so that in particular each $u_{j}\colon\Omega\to\mathbb{C}$ is $K$-quasiregular and, in view of the Dirichlet condition, the sequence $(u_{j})$ is bounded in $\mathrm{W}^{1,2}(\Omega)$. By a normality result for $K$-quasiregular maps [5], $(u_{j})$ admits a subsequence $(u_{j_{k}})$ that converges uniformly to a $K$-quasiregular map $u\in g+\mathrm{W}^{1,2}_{0}(\Omega)$. In order to conclude that $u$ is a minimizer one therefore needs lower semicontinuity of $\mathcal{B}[v]$ along the uniformly convergent subsequence $(u_{j_{k}})$. Apart from the technical assumption on the limit map $u$ the following result yields exactly that.
Theorem 1.5.
Let $p>2$, $K\equiv p/(p-2)$, $\Omega$ be an open subset of $\mathbb{C}$ and suppose $u_{j}$, $u\colon\Omega\to\mathbb{C}$ are $K$-quasiregular maps such that $u_{j}\to u$ locally uniformly on $\Omega$. If $u\in\mathrm{W}^{1,p}_{\operatorname{loc}}(\Omega)$, then
(1.9) $\int_{\Omega^{\prime}}\!\mathbf{B}_{p}(\textup{D}u)\,\mathrm{d}m(z)\leq\liminf_{j\to\infty}\int_{\Omega^{\prime}}\!\mathbf{B}_{p}(\textup{D}u_{j})\,\mathrm{d}m(z)$
holds for each subset $\Omega^{\prime}\Subset\Omega$ with $\mathscr{L}^{2}(\partial\Omega^{\prime})=0$.
We refer to Corollary 3.7 for a global version of the Burkholder lower semicontinuity result on Dirichlet classes. While the result is stated for the Burkholder functional the interested reader will see that the proof also applies to the principal quasiconvex functionals of Theorem 2.2.
Note that in Theorem 1.5, since we do not know if the sequence $\mathbf{B}_{p}(\textup{D}u_{j})$ is locally equi-integrable, we need to handle the concentration and oscillation phenomena simultaneously. In this setting we will actually prove that the singular part of a suitably defined concentration measure (the reduced defect measure) is zero.
We next consider lower semicontinuity results for general principal quasiconvex functionals, which are consistent with the blow up condition
(1.10) ${\bf E}(A)\textrm{ diverges as }\det A\to 0^{+}.$
The growth conditions we are able to deal with are the following,
(1.11) ${\bf E}^{-}(A)\leqslant C_{1}(\max\{|A|^{2-\epsilon},-\log(\det A),K_{A}\}+C_{2},$
where ${\bf E}^{-}(A)=\left|\min\{{\bf E}(A),0\}\right|$. With these bounds we have:
Theorem 1.6.
Let ${\bf E}\colon\mathbb{R}^{2\times 2}\to\mathbb{R}\cup\{+\infty\}$ be lower semicontinuous, $\mathrm{W}^{1,1}$-principal quasiconvex and satisfying (1.11). Fix a homeomorphism $g\colon\mathbb{C}\to\mathbb{C}$ of class $\mathrm{W}^{1,2}_{\operatorname{loc}}(\mathbb{C})$ and assume it has integrable distortion on $\Omega$, $\|K_{g}\|_{\mathrm{L}^{1}(\Omega)}<+\infty$. Let $(u_{j})\subset g+\mathrm{W}^{1,2}_{0}(\Omega)$ with $\sup_{j}\|K_{u_{j}}\|_{\mathrm{L}^{q}(\Omega)}<\infty$ for some $q>1$ and suppose that $u_{j}\rightharpoonup u$ in $\mathrm{W}^{1,2}(\Omega)$, Then
$\liminf_{j\to\infty}\int_{\Omega}{\bf E}(\textup{D}u_{j}(z))\,\mathrm{d}m(z)\geq\int_{\Omega}{\bf E}(\textup{D}u(z))\,\mathrm{d}m(z).$
Our last result provides lower semicontinuity along sequences of maps that have asymptotically finite distortion.
Definition 1.7.
Let $(u_{j})$ be a sequence in $\mathrm{W}^{1,2}_{\operatorname{loc}}(\Omega)$ and $\mathbb{A}$ be a measurable subset of $\Omega$. We say that the maps have asymptotically finite distortion on $\mathbb{A}$ if
(1.12) $K(x)\equiv\limsup_{j\to\infty}\mathbf{K}(\textup{D}u_{j}(x))<+\infty\quad\mbox{ for a.e.}\;x\in\mathbb{A}.$
Our next result is a general lower semicontinuity result for sequences with asymptotically finite distortion. We emphasize that the mappings do not necessarily have integrable distortion and that the result can be localized at an arbitrary measurable set $\mathbb{A}$.
Theorem 1.8.
Let $\mathbf{E}\colon\mathbb{R}^{2\times 2}\to\mathbb{R}\cup\{+\infty\}$ be Borel, lower semicontinuous on +$\mathbb{R}^{2\times 2}_{+}\cup\{0\}$, $\mathrm{W}^{1,2}$-principal quasiconvex and assume that
(1.13) $\bigl{|}\mathbf{E}(A)\bigr{|}\leq\mathbf{C}(K_{A})\bigl{(}|A|^{2}+1\bigr{)},\quad\forall A\in\mathbb{R}^{2\times 2}_{+}\cup\{0\}.$
Let $\mathbb{A}$ be a measurable subset of $\Omega$.
Then, if $\,(u_{j})\,\subset\mathrm{W}^{1,2}_{\operatorname{loc}}(\Omega)$ has asymptotically finite distortion on $\mathbb{A}$, if the sequence converges weakly to $u$ in $\mathrm{W}^{1,2}_{\operatorname{loc}}(\Omega)$ and if
(1.14) $\bigl{(}\mathbf{E}(\textup{D}u_{j})^{-}\bigr{)}\,\mbox{ is equi-integrable on }\mathbb{A},$
then we have
$\liminf_{j\to\infty}\int_{\mathbb{A}}\!\mathbf{E}(\textup{D}u_{j})\,\mathrm{d}m(z)\geq\int_{\mathbb{A}}\!\mathbf{E}(\textup{D}u)\,\mathrm{d}m(z).$
Remark 1.9.
Notice that equi-integrability of the negative part in (1.14) is needed even in the case of the determinant.
In this introduction we have stated the homogeneous versions of the lower semicontinuity estimates, but in the bulk of the text we will deal with general non autonomous $z$-dependent versions of the integrands. Before we embark on the proofs we present examples of principal quasiconvex functionals in section 2. In section 3 we describe the proofs. The proofs of Theorems 1.8 and 1.6 are of the same nature. First we use localization, we then deal with the gradient Young measures via Stoilow factorization, and finally we apply principal quasiconvexity. The localization for the case of asymptotic finite distortion is however rather delicate, as for example the 0-1 law from [3] is used. The proof of Theorem 1.5 is more subtle as we cannot rely on gradient Young measure theory. Instead we analyze the limit measure of the $\mathrm{L}^{1}$ bounded sequence $\mathbf{B}_{p}(\textup{D}u_{j})$ and localize differently at points where the limit measure is absolutely continuous with respect to Lebesgue measure or singular. This blow up technique adjusts very well to principal quasiconvexity, though to treat the singular part we need to assume critical integrability for the limit map (not for the sequence though).
Acknowledgments
D.F, K.A, A.K acknowledge the financial support of QUAMAP, the ERC Advanced Grant 834728, and of the Severo Ochoa Programme CEX2019-000904-S. A.G. was supported by Dr. Max Rössler, the Walter Haefner Foundation and the ETH Zürich Foundation. D.F and A.K were partially supported by CM and UAM, and A.K by Academy of Finland CoE Randomness and Structures, and Academy Fellowship Grant 355840. D.F acknowledges financial support by PI2021-124-195NB-C32. D.F also wants to acknowledge the hospitality of the Mathematical Institute of Oxford during the summer of 2023.
2. Examples of Principal quasiconvex functionals
In this section we give examples of functionals satisfying the condition of Definition 1.2. A first simple class of examples is furnished by the following result:
Proposition 2.1.
Assume $\Phi\colon\mathbb{R}^{2\times 2}\times\mathbb{R}\to\mathbb{R}\cup\{+\infty\}$ is convex and that the partial function $t\mapsto\Phi(A,t)$ is non-increasing for each fixed $A\in\mathbb{R}^{2\times 2}$. Then the functional $\mathbf{F}(A)\equiv\Phi(A,\det A)$, $A\in\mathbb{R}^{2\times 2}_{+}$, is $\mathrm{W}^{1,2}$-principal quasiconvex.
The proof follows easily if we combine Jensen’s inequality for convex functions with the area formula for principal maps of class $\mathrm{W}^{1,2}_{\operatorname{loc}}(\mathbb{C})$ (see [5, Th. 2.10.1]). The details are left to the interested reader and we only remark here that a prime example of a polyconvex functional that is also $\mathrm{W}^{1,2}$-principal quasiconvex is the distortion functional $\mathbf{K}(A)\equiv K_{A}\equiv|A|^{2}/\det A$ for $\det A>0$.
We now turn to the example from which all our research emanates, namely the well-known Burkholder functional $\mathbf{B}_{p}\colon\mathbb{R}^{2\times 2}\to\mathbb{R}$ that we consider for exponents $p\geq 2$. Recall that in complex notation it is given as
(2.1) $\mathbf{B}_{p}(A)=\bigl{(}(p-1)|a_{-}|-|a_{+}|\bigr{)}\bigl{(}|a_{+}|+|a_{-}|\bigr{)}^{p-1}$
for a matrix $A\in\mathbb{R}^{2\times 2}$ that has conformal-anticonformal coordinates $(a_{+},a_{-})\in\mathbb{C}^{2}$. It is arguably the prime example of a rank-one convex functional candidate to be quasiconvex; in general, rank-one convexity is not sufficient for quasiconvexity [45]. On one hand, the Burkholder functional is structurally simple, since it is isotropic (that is, invariant under pre- and post-composition with rotations) and $p$-homogeneous. On the other, it is tightly linked to the optimal higher integrability properties of quasiconformal mappings, see [2]. In the recent work, [4] it was shown that the Burkholder functional is quasiconvex when restricted to quasiconformal test functions with appropriate distortion. In other words, the local Burkholder functional $\mathbf{B}_{K}$ defined in the introduction is quasiconvex. As described in [4], when taking the limit $p$ going to $2$ this quasiconvexity has interesting consequences also for existence of minimizers in elasticity.
In [4] it was proven that both $W(A)=K_{A}-\log K_{A}+\log(\det A)$ and the local Burkholder functional are principal quasiconvex. Our next result pinpoints which are the properties of the Burkholder functional which lead to its principal quasiconvexity:
Theorem 2.2.
Let $p>2$ and $\mathbf{F}\colon\mathbb{R}^{2\times 2}\to\mathbb{R}$ be a $p$-homogeneous, isotropic and quasiconvex functional on $Q_{2}(K)$ with $K\equiv p/(p-2)$. The quasiconvexity condition required is more precisely that
[…]
holds for all $A\in Q_{2}(K)$ and $\phi\in\mathrm{W}^{1,2}_{0}(\mathbb{D})$ with $A+\textup{D}\phi(z)\in Q_{2}(K)$ a.e.
Assume $\mathbf{F}(A)\leq 0$ for all $A\in Q_{2}(K)$ and define the local functional
$\mathbf{F}_{K}(A)\equiv\left\{\begin{array}[]{ll}\mathbf{F}(A)&\mbox{ if }A\in Q_{2}(K)\\ +\infty&\mbox{ if }A\in\mathbb{R}^{2\times 2}\setminus Q_{2}(K).\end{array}\right.$
Then $\mathbf{F}_{K}$ is $\mathrm{W}^{1,2}$-principal quasiconvex.
The proof is exactly the same as the proof in [4], as under the assumed symmetries it holds that for $\alpha\in\mathbb{C}$ and $A\in\mathbb{R}^{2\times 2}$, $\mathbf{F}(\alpha A)=|\alpha|^{p}|\mathbf{F}(A)$ which is the key additional property of $\mathbf{B}_{p}$ used in [4]. In fact, as for the local Burkholder functional, also $\mathbf{F}_{K}$ is actually $\mathrm{W}^{1,\frac{2K}{K+1}}$-principal quasiconvex: this is so because a weakly $K$-quasiregular map in this class is in $\mathrm{W}^{1,2}_{\operatorname{loc}}$ (see [5] and the references therein).
We now give a larger, more general family of examples, with different symmetries. In order to motivate these examples we recall that the profound implications of the restricted quasiconvexity of the Burkholder functional on the higher integrability theory of quasiconformal maps led the authors of [8] to define and investigate a more general family of functionals, so called complex Burkholder functionals. These functionals describe not only the integrability properties of quasiconformal mappings, loosely speaking corresponding to the stretch of such maps, but also the rotation of such maps. The definition is a bit cumbersome but very natural from the quasiconformal view point, and we now describe it in detail.
It was discovered in [6, 8] that a fruitful way to understand the Burkholder functional is to see $\int_{\mathbb{D}}\mathbf{B}_{p}(\textup{D}f)\,\mathrm{d}m(z)$ as a weighted $\mathrm{L}^{p}$ norm of $\textup{D}f$, where the weight depends on $p$ and $f$: indeed, first recall that any $K$-quasiconformal map is a $\mathrm{W}^{1,2}_{\operatorname{loc}}$-solution to the Beltrami equation
$f_{\overline{z}}=\mu(z)f_{z},$
where $\mu=\mu_{f}$ satisfies $|\mu_{f}(z)|\leq\frac{K-1}{K+1}<1$ for a.e. $z$. Further, note that for such a map we may write
$\mathbf{B}_{p}(\textup{D}f)=\Big{(}\frac{p|\mu_{f}|}{1+|\mu_{f}|}-1\Big{)}\Bigl{|}(1+|\mu_{f}|)f_{z}\Bigr{|}^{p}.$
In order to generalize this definition to describe the optimal complex powers for quasiconformal maps we take the exponent $p$ to be a complex number which satisfies
(2.2) $1\leq|\mathtt{p}-1|,\qquad\mathtt{p}\neq 0,$
and we define $\beta=\beta(\mathtt{p})$ as the complex number determined uniquely through the equations
$|\beta|+|\beta-2|=2|\mathtt{p}-1|\quad\text{ and }\quad\operatorname{Re}(\beta/\mathtt{p})=1.$
Then we set
$\mathbf{B}_{p}(A)=\Big{(}\frac{\mathtt{p}\eta|\mu_{A}|}{1+\eta|\mu_{A}|}-1\Big{)}|(a_{+}+\eta|\mu_{A}|a_{+})^{\beta}|.$
Here $\eta$ depends on $|\mu_{A}|$ and $\mathtt{p}$ through the implicit relations
$|\eta(z)|=1\quad\text{ and }\quad\textup{arg}(p\eta)=\textup{arg}(1+\eta|\mu_{A}|).$
We readily see that the structure of the complex Burkholder functional is similar to that of $B_{p}$ but it contains complex powers of $f_{z}$. Thus, as mentioned above, its quasiconvexity relates not only to integrabilty issues but also to how fast quasiconformal maps can wind a line into a spiral. From the view of the structure of the functional, for a complex parameter $\beta$, the functional is $(\operatorname{Re}\beta)$-homogeneous and the group of symmetries is described by the logarithmic spiral $S_{\beta}=\{z:|z^{\beta}|=1\}$. We invite the reader to find a function $b\colon\mathbb{R}\to\mathbb{C}$ such that $\mathbf{B}_{p}(A)=|a_{+}^{\beta}|b(|\mu_{A}|)$. In a companion paper we prove that the full family of local complex Burkholder functionals is principal quasiconvex:
Theorem 2.3.
Let $\mathbf{B}_{\mathtt{p}}$ be the $\mathtt{p}$-Burkholder functional with $\mathtt{p}$ satisfying condition (2.2). Then $\bigl{\{}\mathbf{B}_{\mathtt{p}}\leq 0\bigr{\}}=Q_{2}\left(\frac{|\mathtt{p}-1|+1}{|\mathtt{p}-1|-1}\right)$ and
$\mathcal{B}(A)\equiv\left\{\begin{array}[]{cl}\mathbf{B}_{\mathtt{p}}(A)&\mbox{ if }\mathbf{B}_{\mathtt{p}}(A)\leq 0,\\ +\infty&\mbox{ if }\mathbf{B}_{\mathtt{p}}(A)>0,\end{array}\right.$
is $\mathrm{W}^{1,2}$-principal quasiconvex.
3. Proofs
3.1. Lower semicontinuity along sequences with asymptotically finite distortion
We begin by giving the proof of a more general version of Theorem 1.8. The strategy is as follows: first we apply a general semicontinuity result for Young measures [10]. Then we localize and show that our assumed growth conditions and principal quasiconvexity imply that the integrands satisfy the Jensen inequality for the corresponding homogeneous gradient Young measures.
Theorem 3.1.
Let $\mathbf{E}\colon\Omega\times\mathbb{R}^{2\times 2}\to\mathbb{R}\cup\{+\infty\}$ be a Borel measurable functional that is +$\infty$+ ∞ on $\Omega\times\bigr{(}\mathbb{R}^{2\times 2}\setminus\mathbb{R}^{2\times 2}_{+}\bigr{)}$ and such that $\mathbf{E}(z,\cdot)$ is lower semicontinuous on +$\mathbb{R}^{2\times 2}_{+}$ for each $z\in\Omega$. Assume that
(3.1) $\bigl{|}\mathbf{E}(z,A)\bigr{|}\leq\mathbf{C}(z,K_{A})\bigl{(}|A|^{2}+1\bigr{)}\quad\forall(z,A)\in\Omega\times\mathbb{R}^{2\times 2}_{+},$
where $\mathbf{C}\colon\Omega\times[1,+\infty)\to[1,+\infty)$ is Borel measurable and $\mathbf{C}(z,\cdot)\colon[1,\infty)\to[1,\infty)$ is an increasing function for each $z\in\Omega$, and that for each $z\in\Omega$, $\mathbf{E}(z,\cdot)$ is $\mathrm{W}^{1,2}$-principal quasiconvex.
Let $(u_{j})$ be a sequence in $\mathrm{W}^{1,2}_{\operatorname{loc}}(\Omega)$ and suppose that the maps have asymptotically finite distortion on a measurable subset $\mathbb{A}$ of $\Omega$:
[…]
Then, if $u_{j}\rightharpoonup u$ in $\mathrm{W}^{1,2}_{\operatorname{loc}}(\Omega)$ and
(3.3) $\bigl{(}\mathbf{E}(\cdot,\textup{D}u_{j})^{-}\bigr{)}\,\mbox{ is equi-integrable on }\mathbb{A},$
we have
$\liminf_{j\to\infty}\int_{\mathbb{A}}\!\mathbf{E}(\cdot,\textup{D}u_{j})\,\mathrm{d}m(z)\geq\int_{\mathbb{A}}\!\mathbf{E}(\cdot,\textup{D}u)\,\mathrm{d}m(z).$
Remark 3.2.
The conclusion of Theorem 3.1 fails in general without the assumption (3.3), but can be reinstated if formulated in terms of biting convergence as in [29]. Under the assumptions of Theorem 3.1 except (3.3) we have that each subsequence of $(u_{j})$ admits a further subsequence $(u_{j_{k}})$ such that
$b\ast\!\lim_{k\to\infty}\mathbf{E}(\cdot,\textup{D}u_{j_{k}})\geq\mathbf{E}(\cdot,\textup{D}u)\quad\mathscr{L}^{2}\mbox{ a.e.~{}in }\mathbb{A}.$
In fact, this is a consequence of Jensen’s inequality for $\mathbf{E}$ and the Young measures generated by subsequences of $(\textup{D}u_{j_{k}})$. That they hold is part of our proof for Theorem 3.1 and we invite the reader to check that their validity do not hinge on the equi-integrability assumption (3.3). Indeed, if $(\nu_{z})_{z\in\Omega}$ is a Young measures generated by a subsequence of $(\textup{D}u_{j_{k}})$, then it is not difficult to see that
$b\ast\!\lim_{k\to\infty}\mathbf{E}(\cdot,\textup{D}u_{j_{k}})\geq\int_{\mathbb{R}^{2\times 2}}\!\mathbf{E}(\cdot,A)\,\mathrm{d}\nu_{\cdot}(A)\geq\mathbf{E}(\cdot,\textup{D}u)\quad\mathscr{L}^{2}\mbox{ a.e.~{}in }\mathbb{A}.$
The key to the proof is the following localization principle for the Young measures that are generated by the derivatives of maps in a weakly converging sequence in $\mathrm{W}^{1,2}_{\operatorname{loc}}$ satisfying the asymptotic distortion condition (3.2). We emphasize that in the subsequent statements and proofs we consider all functions and maps in terms of their precise representatives.
Proposition 3.3.
Let $(u_{j})$ be a bounded sequence in $\mathrm{W}^{1,2}_{\operatorname{loc}}(\Omega)$ of maps satisfying (3.2). If $(\textup{D}u_{j})$ generates the Young measure $(\nu_{z})_{z\in\Omega}$, then for $\mathscr{L}^{2}$ almost all $z\in\mathbb{A}$ the measure $\nu_{z}$ is a homogeneous $\mathrm{W}^{1,2}$ gradient Young measure with support in $Q_{2}\bigl{(}K(z)\bigr{)}$.
Proof.
Put $M_{j}(z)\equiv\sup_{s\geq j}\mathbf{K}(\textup{D}u_{s})$, $z\in\Omega$. Then $M_{j}\colon\Omega\to[1,+\infty]$ is measurable and
(3.4) $M_{j}(z)\geq M_{j+1}(z)\searrow K(z)\,\mbox{ as }\,j\nearrow\infty\,\mbox{ pointwise in }\,z\in\Omega.$
We recall that $K(z)$ is defined at (3.2), where it is also assumed that $K(z)<+\infty$ holds $\mathscr{L}^{2}$ almost everywhere on $\mathbb{A}$. Because $K$ is also clearly measurable it follows from Lusin’s theorem that it is $\mathscr{L}^{2}$ quasi-continuous on $\mathbb{A}$: for each $\varepsilon>0$ there exists a measurable subset $G=G_{\varepsilon}\subset\mathbb{A}$ such that $K$ is finite and continuous relative to $G$ and $\mathscr{L}^{2}(\mathbb{A}\setminus G)<\varepsilon$. Fix an $\varepsilon>0$ and corresponding set $G=G_{\varepsilon}$. Fix $a>1$ and put $G_{j}=G_{j}^{a}\equiv\{z\in G:\,M_{j}(z)<aK(z)\}$ and record that $G_{j}$ are measurable sets that by virtue of (3.4) form an ascending sequence that in the limit $j\to\infty$ exhaust $G$:
(3.5) $G_{j}\subset G_{j+1}\,\mbox{ and }\bigcup_{j\in\mathbb{N}}G_{j}=G.$
We next note that the space $\mathrm{C}_{0}(\mathbb{D}\times\mathbb{R}^{2\times 2})$ with the supremum norm is separable and that we may find countable families $\mathcal{F}_{1}$ in $\mathrm{C}_{c}^{1}(\mathbb{D})$ and $\mathcal{F}_{2}$ in $\mathrm{C}_{c}^{1}(\mathbb{R}^{2\times 2})$ such that the countable family $\mathcal{F}\equiv\bigl{\{}\eta\otimes\Phi:\,\eta\in\mathcal{F}_{1},\,\Phi\in\mathcal{F}_{2}\bigr{\}}$ is total in $\mathrm{C}_{0}(\mathbb{D}\times\mathbb{R}^{2\times 2})$. Consequently, if $(\kappa^{i}_{z})_{z\in\mathbb{D}}$, $i=1,2$, are two Young measures on $\mathbb{D}$ and
$\int_{\mathbb{D}}\!\langle\kappa_{z}^{1},\Phi\rangle\eta(z)\,\mathrm{d}m(z)=\int_{\mathbb{D}}\!\langle\kappa_{z}^{2},\Phi\rangle\eta(z)\,\mathrm{d}m(z)$
holds for all $\eta\otimes\Phi\in\mathcal{F}$, then the Young measures are the same: $\kappa_{z}^{1}=\kappa_{z}^{2}$ for $\mathscr{L}^{2}$ almost all $z\in\mathbb{D}$.
[…]
where $\lambda$ is a positive Radon measure on $\Omega$. In particular we then have for each disk $\mathbb{D}_{r}(z_{0})$ that is compactly contained in $\Omega$ that
$\limsup_{j\to\infty}\mathchoice{{\vbox{\hbox{$\textstyle-$}}\kern-4.86108pt}}{{\vbox{\hbox{$\scriptstyle-$}}\kern-3.25pt}}{{\vbox{\hbox{$\scriptscriptstyle-$}}\kern-2.29166pt}}{{\vbox{\hbox{$\scriptscriptstyle-$}}\kern-1.875pt}}\!\int_{\mathbb{D}_{r}(z_{0})}\!|\textup{D}u_{j}|^{2}\,\mathrm{d}m(z)\leq\frac{\lambda\bigl{(}\overline{\mathbb{D}_{r}(z_{0})}\bigr{)}}{\mathscr{L}^{2}\bigl{(}\mathbb{D}_{r}(z_{0})\bigr{)}},$
and hence by Lebesgue’s differentiation theorem that
(3.7) $\limsup_{r\searrow 0}\limsup_{j\to\infty}\mathchoice{{\vbox{\hbox{$\textstyle-$}}\kern-4.86108pt}}{{\vbox{\hbox{$\scriptstyle-$}}\kern-3.25pt}}{{\vbox{\hbox{$\scriptscriptstyle-$}}\kern-2.29166pt}}{{\vbox{\hbox{$\scriptscriptstyle-$}}\kern-1.875pt}}\!\int_{\mathbb{D}_{r}(z_{0})}\!|\textup{D}u_{j}|^{2}\,\mathrm{d}m(z)\leq\frac{\mathrm{d}\lambda}{\mathrm{d}\mathscr{L}^{2}}(z_{0})<+\infty$
holds for $\mathscr{L}^{2}$ almost all $z_{0}\in\Omega$. We now identify a set of bad points in $\Omega$ that we would like to avoid in the subsequent argument. First recall the countable family $\mathcal{F}_{2}$, and note that for each $\Phi\in\mathcal{F}_{2}$ the function $z\mapsto\langle\nu_{z},\Phi\rangle$ is an $\mathrm{L}^{\infty}(\Omega)$ function, so that $\mathscr{L}^{2}$ almost all $z_{0}\in\Omega$ are Lebesgue points:
$\lim_{r\searrow 0}\mathchoice{{\vbox{\hbox{$\textstyle-$}}\kern-4.86108pt}}{{\vbox{\hbox{$\scriptstyle-$}}\kern-3.25pt}}{{\vbox{\hbox{$\scriptscriptstyle-$}}\kern-2.29166pt}}{{\vbox{\hbox{$\scriptscriptstyle-$}}\kern-1.875pt}}\!\int_{\mathbb{D}_{r}(z_{0})}\!\langle\nu_{z},\Phi\rangle\,\mathrm{d}m(z)=\langle\nu_{z_{0}},\Phi\rangle\in\mathbb{R}.$
Let $N_{\Phi}$ denote the complement in $\Omega$ of this set of Lebesgue points. Next, let $N$ denote the set of points in $\Omega$, where (3.7) fails. Define $\mathcal{N}\equiv N\cup\bigcup_{\Phi\in\mathcal{F}_{2}}N_{\Phi}$ and record that $\mathscr{L}^{2}(\mathcal{N})=0$.
Define for $\mathbb{D}_{r}(z_{0})\Subset\Omega$ the maps
$v_{j,r}(z)\equiv\frac{u_{j}(z_{0}+rz)-(u_{j})_{z_{0},r}}{r},\quad z\in\mathbb{D},$
where $(u_{j})_{z_{0},r}$ is the integral mean of $u_{j}$ over $\mathbb{D}_{r}(z_{0})$. Hereby $v_{j,r}\in\mathrm{W}^{1,2}(\mathbb{D})$ and $(v_{j,r})_{0,1}=0$. Before the next display we introduce some convenient notation for functions $\theta\colon\mathbb{C}\to\mathbb{R}$: the reflected function, $\tilde{\theta}$, is defined as $\tilde{\theta}(z)\equiv\theta(-z)$, and the $\mathrm{L}^{1}$-dilated function with factor $r>0$, $\theta_{r}$, is defined as $\theta_{r}(z)\equiv\theta\bigl{(}z/r\bigr{)}/r^{2}$. In these terms we note that for each $z_{0}\in\Omega\setminus\mathcal{N}$ and $\eta\otimes\Phi\in\mathcal{F}$ the above Lebesgue point property and a standard result about convolution yield (here $\eta$ is extended to $\mathbb{C}\setminus\mathbb{D}$ by 00):
$\lim_{r\searrow 0}\lim_{j\to\infty}\int_{\mathbb{D}}\!\eta\Phi(\textup{D}v_{j,r})\,\mathrm{d}m(z)$ $=$ $\lim_{r\searrow 0}\lim_{j\to\infty}\int_{\Omega}\!\tilde{\eta}_{r}(z_{0}-z)\Phi(\textup{D}u_{j})\,\mathrm{d}m(z)$
$=$ $\lim_{r\searrow 0}\int_{\Omega}\!\tilde{\eta}_{r}(z_{0}-z)\langle\nu_{z},\Phi\rangle\,\mathrm{d}m(z)$
$=$ $\langle\nu_{z_{0}},\Phi\rangle\int_{\mathbb{D}}\!\eta\,\mathrm{d}m(z).$
The bound (3.7) translates to
$\limsup_{r\searrow 0}\limsup_{j\to\infty}\int_{\mathbb{D}}\!|\textup{D}v_{j,r}|^{2}\,\mathrm{d}m(z)\leq\pi\frac{\mathrm{d}\lambda}{\mathrm{d}\mathscr{L}^{2}}(z_{0})<+\infty.$
Fix a point $z_{0}\in G\setminus\mathcal{N}$ that is also a density point for $G$. In addition to the above properties we now also have the following double limit:
$\lim_{r\searrow 0}\lim_{j\to\infty}\frac{\mathscr{L}^{2}\bigl{(}\mathbb{D}_{r}(z_{0})\cap G_{j}\bigr{)}}{\mathscr{L}^{2}\bigl{(}\mathbb{D}_{r}(z_{0})\bigr{)}}=1.$
In order to choose a good null sequence $r_{j}\searrow 0$ we employ a diagonalization argument in connection with the above double limits: Hereby we find for $v_{j}\equiv v_{j,r_{j}}$ that
(3.8) $\lim_{j\to\infty}\int_{\mathbb{D}}\!\eta\Phi(\textup{D}v_{j})\,\mathrm{d}m(z)=\int_{\mathbb{D}}\!\eta\,\mathrm{d}m(z)\langle\nu_{z_{0}},\Phi\rangle\quad\forall\,\eta\otimes\Phi\in\mathcal{F},$
(3.9) $\sup_{j\in\mathbb{N}}\int_{\mathbb{D}}\!|\textup{D}v_{j}|^{2}\,\mathrm{d}m(z)<+\infty$
and
(3.10) $\lim_{j\to\infty}\frac{\mathscr{L}^{2}\bigl{(}\mathbb{D}_{r_{j}}(z_{0})\cap G_{j}\bigr{)}}{\mathscr{L}^{2}\bigl{(}\mathbb{D}_{r_{j}}(z_{0})\bigr{)}}=1.$
Because $(v_{j})_{0,1}=0$ we deduce using Poincaré’s inequality and (3.9) that the sequence $(v_{j})$ is bounded in $\mathrm{W}^{1,2}(\mathbb{D})$. Therefore any subsequence of $(\textup{D}v_{j})$ admits a further subsequence that generates a Young measure. By (3.8) this Young measure is the homogeneous Young measure $(\nu_{z_{0}})_{z\in\mathbb{D}}$. A routine argument then shows that the full sequence $(\textup{D}v_{j})$ generates the homogeneous Young measure $(\nu_{z_{0}})_{z\in\mathbb{D}}$. It is then easy to check that $(v_{j})$ converges weakly in $\mathrm{W}^{1,2}(\mathbb{D})$ to the linear map $\langle\nu_{z_{0}},\mathrm{Id}\rangle$. In particular it follows that $(\nu_{z_{0}})_{z\in\mathbb{D}}$ is a homogeneous $\mathrm{W}^{1,2}$ gradient Young measure. We use (3.10) to get information about the support of $\nu_{z_{0}}$. Recall the parameter $a>1$ used in the definition of $G_{j}$ and put
$\mathcal{E}=\mathcal{E}_{a}\equiv\bigl{\{}A\in\mathbb{R}^{2\times 2}:\,|A|^{2}>a^{2}K(z_{0})\det A\bigr{\}}.$
This is clearly an open set in $\mathbb{R}^{2\times 2}$ and we note that for $z\in\mathbb{D}$ we have $\textup{D}v_{j}(z)\in\mathcal{E}$ precisely when $|\textup{D}u_{j}(z_{0}+r_{j}z)|^{2}>a^{2}K(z_{0})\det\textup{D}u_{j}(z_{0}+r_{j}z)$. Now recall that $K$ is continuous and finite relative to $G$ and so for some $j_{a}\in\mathbb{N}$ we have $K(z_{0}+r_{j}z)<aK(z_{0})$ for all $z\in\mathbb{D}\cap\bigl{(}(G-z_{0})/r_{j}\bigr{)}$ for $j\geq j_{a}$. In view of the definition of $G_{j}$ we therefore have when $j\geq j_{a}$ that
(3.11) $|\textup{D}u_{j}(z_{0}+r_{j}z)|^{2}$$\leq$ $aK(z_{0}+r_{j}z)\det\textup{D}u_{j}(z_{0}+r_{j}z)$
$\leq$ $a^{2}K(z_{0})\det\textup{D}u_{j}(z_{0}+r_{j}z)$
holds for all $z\in\mathbb{D}\cap\bigl{(}(G_{j}-z_{0})/r_{j}\bigr{)}$. Because $\mathcal{E}$ is an open set its indicator function $\mathbf{1}_{\mathcal{E}}$ is lower semicontinuous, so we can find functions $\Phi_{k}\in\mathrm{C}_{c}(\mathbb{R}^{2\times 2})$ such that $0\leq\Phi_{k}\leq\Phi_{k+1}\nearrow\mathbf{1}_{\mathcal{E}}$ as $k\nearrow+\infty$. Now for each $k\in\mathbb{N}$,
$\liminf_{j\to\infty}\mathchoice{{\vbox{\hbox{$\textstyle-$}}\kern-4.86108pt}}{{\vbox{\hbox{$\scriptstyle-$}}\kern-3.25pt}}{{\vbox{\hbox{$\scriptscriptstyle-$}}\kern-2.29166pt}}{{\vbox{\hbox{$\scriptscriptstyle-$}}\kern-1.875pt}}\!\int_{\mathbb{D}}\!\mathbf{1}_{\mathcal{E}}(\textup{D}v_{j})\,\mathrm{d}m(z)\geq\liminf_{j\to\infty}\mathchoice{{\vbox{\hbox{$\textstyle-$}}\kern-4.86108pt}}{{\vbox{\hbox{$\scriptstyle-$}}\kern-3.25pt}}{{\vbox{\hbox{$\scriptscriptstyle-$}}\kern-2.29166pt}}{{\vbox{\hbox{$\scriptscriptstyle-$}}\kern-1.875pt}}\!\int_{\mathbb{D}}\!\Phi_{k}(\textup{D}v_{j})\,\mathrm{d}m(z)=\langle\nu_{z_{0}},\Phi_{k}\rangle$
and so taking $k\nearrow+\infty$ we get by virtue of Lebesgue’s monotone convergence theorem
$\liminf_{j\to\infty}\mathchoice{{\vbox{\hbox{$\textstyle-$}}\kern-4.86108pt}}{{\vbox{\hbox{$\scriptstyle-$}}\kern-3.25pt}}{{\vbox{\hbox{$\scriptscriptstyle-$}}\kern-2.29166pt}}{{\vbox{\hbox{$\scriptscriptstyle-$}}\kern-1.875pt}}\!\int_{\mathbb{D}}\!\mathbf{1}_{\mathcal{E}}(\textup{D}v_{j})\,\mathrm{d}m(z)\geq\nu_{z_{0}}\bigl{(}\mathcal{E}\bigr{)}.$
To conclude the proof we estimate the right-hand side using (3.11) for $j\geq j_{a}$:
[…]
$\leq$ $\frac{\mathscr{L}^{2}\bigl{(}\mathbb{D}_{r_{j}}(z_{0})\setminus G_{j}\bigr{)}}{\mathscr{L}^{2}\bigl{(}\mathbb{D}_{r_{j}}(z_{0})\bigr{)}}$
and the latter tends by (3.10) to 00 as $j\to\infty$. Consequently, $\nu_{z_{0}}\bigl{(}\mathcal{E}\bigr{)}=0$. Recall that $\mathcal{E}=\mathcal{E}_{a}$ and that $a>1$ is arbitrary, so by monotone convergence we deduce that $\nu_{z_{0}}\bigl{(}\bigcup_{a>1}\mathcal{E}_{a}\bigr{)}=0$. But then $\nu_{z_{0}}$ is carried by $Q_{2}(K(z_{0}))$, and since $Q_{2}(K(z_{0}))$ is a closed set it must contain the support of $\nu_{z_{0}}$, as required. The above holds for all $z_{0}\in G\setminus\mathcal{N}$ that are density points of $G$, where we recall that $G$ was a measurable subset of $\mathbb{A}$ such that $\mathscr{L}^{2}(\mathbb{A}\setminus G)<\varepsilon$. Here $\varepsilon>0$ was arbitrary and it is then routine to conclude the proof. ∎
Proof of Theorem 3.1..
Because of (3.3) we can without loss of generality assume, considering a suitable subsequence that we for notational convenience do not relabel, that
$\lim_{j\to\infty}\int_{\mathbb{A}}\mathbf{E}(\cdot,\textup{D}u_{j})\,\mathrm{d}m(z)=\liminf_{j\to\infty}\int_{\mathbb{A}}\mathbf{E}(\cdot,\textup{D}u_{j})\,\mathrm{d}m(z)\in\mathbb{R}$
and that the sequence $(\textup{D}u_{j})$ generates the Young measure $(\nu_{z})_{z\in\Omega}$. Here, for each $z\in\Omega$, the partial functional $\mathbf{E}(z,\cdot)$ is lower semicontinuous on +$\mathbb{R}^{2\times 2}_{+}$, and so it coincides with its lower semicontinuous envelope
$\mathbf{E}_{\ast}(z,A)\equiv\liminf_{A^{\prime}\to A}\mathbf{E}(z,A^{\prime})$
there. Note that $\mathbf{E}_{\ast}\colon\Omega\times\mathbb{R}^{2\times 2}\to[-\infty,+\infty]$ is Borel measurable and that
$\lim_{j\to\infty}\int_{\Omega}\mathbf{E}(\cdot,\textup{D}u_{j})\,\mathrm{d}m(z)=\lim_{j\to\infty}\int_{\Omega}\mathbf{E}_{\ast}(\cdot,\textup{D}u_{j})\,\mathrm{d}m(z)\in\mathbb{R},$
where again the negative parts $\bigl{(}\mathbf{E}_{\ast}(\cdot,\textup{D}u_{j})^{-}\bigr{)}$ form an equi-integrable sequence on $\mathbb{A}$. We may then appeal to a general lower semicontinuity result from Young measure theory [10] whereby
$\lim_{j\to\infty}\int_{\mathbb{A}}\mathbf{E}_{\ast}(\cdot,\textup{D}u_{j})\,\mathrm{d}m(z)\geq\int_{\mathbb{A}}^{\ast}\!\langle\nu_{z},\mathbf{E}_{\ast}(z,\cdot)\rangle\,\mathrm{d}m(z).$
It is at this stage we employ the localization principle from Proposition 3.3. Accordingly $\nu_{z}$ is, for $\mathscr{L}^{2}$ almost all $z\in\mathbb{A}$, a homogeneous $\mathrm{W}^{1,2}$ gradient Young measure supported on $Q_{2}\bigl{(}K(z)\bigr{)}$. Because $\mathbf{E}(z,\cdot)$ in particular is lower semicontinuous there we infer that
$\langle\nu_{z},\mathbf{E}_{\ast}(z,\cdot)\rangle=\langle\nu_{z},\mathbf{E}(z,\cdot)\rangle\quad\mbox{ for $\mathscr{L}^{2}$ a.e. }z\in\mathbb{A}.$
Moreover, $\mathbf{E}(z,\cdot)$ is principal quasiconvex and real-valued on +$\mathbb{R}^{2\times 2}_{+}$, so must be rank-one convex and thus locally Lipschitz on +$\mathbb{R}^{2\times 2}_{+}\setminus\{0\}$. It need not be continuous at 00 relative to +$\mathbb{R}^{2\times 2}_{+}$, but is lower semicontinuous there by assumption.
Next, for $\mathscr{L}^{2}$ a.e. $z\in\mathbb{A}$, from [3] we find a sequence of principal $K(z)$-quasiconformal maps $f_{j}\colon\mathbb{C}\to\mathbb{C}$ such that $f_{j}\rightharpoonup\langle\nu_{z},\mathrm{Id}\rangle$ in $\mathrm{W}^{1,p}(\mathbb{D})$ for each $p<2K(z)/(K(z)-1)$ and $(\textup{D}f_{j})$ generates $\nu_{z}$. Recall that according to the 0-1 Law for homogeneous gradient Young measures from [3] we have that either $\nu_{z}(\{0\})=0$ or $\nu_{z}=\delta_{0}$. For measures $\nu_{z}$ of the former type, $\mathbf{E}(z,\cdot)$ is therefore continuous $\nu_{z}$ almost everywhere and so as the quadratic growth (3.1) ensures that the sequence $\bigl{(}\mathbf{E}(\textup{D}f_{j})\bigr{)}$ is equiintegrable on $\mathbb{D}$, we arrive at
[…]
Since we also have that $f_{j}\rightharpoonup A_{f}$ in $\mathrm{W}^{1,2}(\mathbb{D})$, where $\ A_{f}=\langle\nu_{z},\mathrm{Id}\rangle$, we get that $A_{f_{j}}\to A_{f}$. It now follows directly from principal quasiconvexity that $\langle\nu_{z},\mathbf{E}(z,\cdot)\rangle\geq\mathbf{E}\bigl{(}z,\langle\nu_{z},\mathrm{Id}\rangle\bigr{)}$ and this concludes the proof for measures with $\nu_{z}(\{0\})=0$.
For the other measures $\nu_{z}=\delta_{0}$ we simply use that $\mathbf{E}(z,\cdot)$ is lower semicontinuous at 00, that the quadratic growth (3.1) still ensures that the sequence $\bigl{(}\mathbf{E}(\textup{D}f_{j})\bigr{)}$ is equiintegrable on $\mathbb{D}$ (in particular the negative parts) and then the general lower semicontinuity result for Young measures [10]. This concludes the proof. ∎
3.2. Lower semicontinuity with integrable distortion
In this subsection we give the proof of a more general version of Theorem 1.6, which follows the proof in [4, Section 12] concerning the lower semicontinuity of the non polyconvex functional $W$. This functional was introduced in [6] and was also studied in the recent works [46, 47]. As the result below shows, the crucial feature of $W$ is its principal quasiconvexity.
Theorem 3.4.
Let $\mathbf{E}\colon\Omega\times\mathbb{R}^{2\times 2}\to\mathbb{R}\cup\{+\infty\}$ be a Borel measurable functional that is +$\infty$+ ∞ on $\Omega\times\bigr{(}\mathbb{R}^{2\times 2}\setminus\mathbb{R}^{2\times 2}_{+}\bigr{)}$ and such that $\mathbf{E}(z,\cdot)$ is lower semicontinuous on +$\mathbb{R}^{2\times 2}_{+}$ for each $z\in\Omega$ and principal quasiconvex. Assume that there exists constants $C_{1},C_{2}$ such that
(3.12) ${\bf E}^{-}(z,A)\leqslant C_{1}(\max\{|A|^{2-\epsilon},-\log(\det A),K_{A}\}+C_{2}$
Let $g$ be a globally defined homeomorphism in $\mathbb{C}$ with $K_{g}\in L^{1}$ and consider a sequence of maps $u_{j}\in g+\mathrm{W}^{1,2}_{0}(\Omega)$ with $K_{u_{j}}\in L^{q}$ for $q>1$. Then, if $u_{j}\rightharpoonup u$ in $\mathrm{W}^{1,2}_{\operatorname{loc}}(\Omega)$ we have
$\liminf_{j\to\infty}\int_{\Omega}\!\mathbf{E}(\cdot,\textup{D}u_{j})\,\mathrm{d}m(z)\geq\int_{\Omega}\!\mathbf{E}(\cdot,\textup{D}u)\,\mathrm{d}m(z).$
The proof follows the rough strategy used in the previous subsection. As a first step towards this theorem, we first show that the growth conditions assumed are enough for equiintegrability: this relies on [33, 32] as presented in [4, Proposition 12.8]. By Young measure theory the proof then reduces to obtaining Jensen’s inequality for a homogeneous gradient Young measure, to which Theorem 1.4 then applies. The conclusion then follows from principal quasiconvexity.
Proposition 3.5.
Let ${\bf E}$ satisfy (1.11) and let $(u_{j})\subset\mathrm{W}^{1,2}_{g}(\Omega)$ be a sequence bounded in $\mathrm{W}^{1,2}(\Omega)$, and $\sup_{j}\|K_{u_{j}}\|_{\mathrm{L}^{q}(\Omega)}<\infty$. Then ${\bf E}(\textup{D}u_{j}))$ is equiintegrable in $\Omega$.
Proof.
Recall (1.11) and that by our assumptions both $|\textup{D}u_{j}|^{2-\epsilon}$ and $K_{u_{j}}$ are equiintegrable. It remains to show that $\log J_{u_{j}}$ is equiintegrable but this follows from the pointwise estimate:
$|\log(J_{u_{j}})|\leq\log\left(e+\frac{1}{J_{u_{j}}}\right)+(J_{u_{j}})^{\frac{1}{2}}.$
Namely the first term is directly taken care by [4, Proposition 12.8], and the second term from the boundary values. ∎
Proof of theorem 3.4.
First note that, by the boundary values, the sequence $u_{j}$ consists of homeomorphisms. By passing to subsequences we may assume that $u_{j}\rightharpoonup u$ in $\mathrm{W}^{1,2}(\Omega)$ and, by the fundamental theorem of Young Measures [4, Theorem 2.8], that $u_{j}$ generates the $\mathrm{W}^{1,2}$-gradient Young measure $(\nu_{z})_{z\in\Omega}$. Since $({\bf E}(z,\textup{D}u_{j}))$ is equi-integrable by Proposition 3.5, we have
(3.13) $\lim_{j\to\infty}\int_{\Omega}{\bf E}(z,\textup{D}u_{j})\,\mathrm{d}m(z)=\int_{\Omega}\int_{\mathbb{R}^{2\times 2}}{\bf E}(z,A)\,\mathrm{d}\nu_{z}(A)\,\mathrm{d}m(z).$
For a.e. $z\in\Omega$ we have
(3.14) $\det\overline{\nu}_{z}>0\quad\textup{and}\quad\nu_{z}\in\mathscr{M}^{2}_{\textup{qc}},$
where the first claim follows by considering the functional $\min\{0,\det\}$ and applying it to the Young measure, and the latter claim due to Proposition 3.3.
Let us fix a point $z\in\Omega$ for which (3.14) holds. The measure $\nu_{z}$ is generated by taking a diagonal subsequence $(\psi_{j})$ of $\psi_{j,\lambda}(w)=\lambda^{-1}u_{j}(z_{0}+\lambda w)$, where $j\to\infty$ and $\lambda\to 0$, cf.  [4, Theorem 2.8]. In particular, $\psi_{j}\colon\mathbb{D}\to\mathbb{C}$ is a sequence of homeomorphisms such that $\sup_{j}\|K_{\psi_{j}}\|_{\mathrm{L}^{q}(\mathbb{D})}<\infty$. Then we are precisely in position to apply to this sequence Theorem 1.4 to find sequences $(h_{j}),(f_{j})$ of respectively conformal and principal maps. Let us call $A_{z}$ the center of mass of $\nu_{z}$, and notice that $f_{j}\to A_{z}$ locally uniformly in $\mathbb{C}$, $\textup{D}f_{j}$ converges weakly to $A_{z}$, $A_{f_{j}}=\int_{\textup{D}}\textup{D}f_{j}\,\mathrm{d}m(z)$ converges to $A_{z}$, and $K_{f_{j}}=K_{\psi_{j}}$ a.e. in $\mathbb{D}$. Therefore as in Proposition 3.5 it follows that ${\bf{E}}(z,\textup{D}f_{j}(w))$ equi-integrable over $\mathbb{D}$. Notice that $f_{j}$ is globally defined and $K_{f_{j}}\equiv 1$ off the unit disc. Hence the proof of Jensen inequality finishes as follows:
$\int_{\mathbb{R}^{2\times 2}}{\bf E}(z,A)\,\mathrm{d}\nu_{z}(A)$ $=\lim_{j\to\infty}\fint_{\mathbb{D}(0,R)}{\bf E}(z,\textup{D}f_{j}(w))\,\mathrm{d}m(w)$
$\to\lim_{j\to\infty}\fint_{\mathbb{D}}{\bf E}(z,,\textup{D}f_{j}(w))\,\mathrm{d}m(w)\geqslant\lim_{j\to\infty}{\bf E}(A_{f_{j}})=E(A_{z})$
as $R\nearrow 1$, Inserting the Jensen inequality for each $\nu_{z}$ in (3.13) yields the required lower semicontinuity. ∎
3.3. Lower Semicontinuity of the critical Burkholder energy
3.3.1. An integrability property of the Burkholder functional
Before starting the proof we record the following integrability property of the Burkholder functional. That a result of this type holds is well-known [5], but our statement here is more explicit.
Proposition 3.6 ($\mathrm{L}^{1}$ boundedness).
Let $p>2$ and put $K\equiv p/(p-2)$. There exists a constant $C=C(p)$ with the following property. If $u\colon\Omega\to\mathbb{C}$ is a $K$-quasiregular map and $\mathbb{D}_{R}(z_{0})\Subset\Omega$, then for all $\delta\in(0,1)$ we have
[…]
The point here is, as was mentioned in the Introduction, that a $K$-quasiregular map need not be of class $\mathrm{W}^{1,p}_{\operatorname{loc}}$. That the Burkholder functional is locally integrable despite being nonpositive and $p$-homogeneous on the derivative of a $K$-quasiregular map comes down to it vanishing on the boundary of the $K$-quasiconformal well, where the large values of the derivative must concentrate.
Proof.
Changing variables we may without loss of generality assume that $z_{0}=0$ and $R=1$. We then Stoilow factorize $u$ on $\mathbb{D}$, whereby $u=h\circ f$ for a $K$-quasiconformal principal map $f\colon\mathbb{C}\to\mathbb{C}$ and a holomorphic map $h\colon f(\mathbb{D})\to\mathbb{C}$. If $m\equiv\sup_{\mathbb{D}}|u|$, then also $m=\sup_{f(\mathbb{D})}|h|$ and so by Cauchy’s integral formula and Hölder properties of $K$-quasiconformal principal maps [5] we estimate
$\sup_{\delta\mathbb{D}}\bigl{|}h^{\prime}\circ f\bigr{|}\leq\frac{m}{\operatorname{dist}(f(\delta\mathbb{D}),\partial f(\mathbb{D}))}\leq c_{p}\frac{m}{(1-\delta)^{K}}.$
Next, using that $\mathbf{B}_{p}$ is nonpositive and the Burkholder area inequality [4]:
$\int_{\delta\mathbb{D}}\!\mathbf{B}_{p}(\textup{D}u)\,\mathrm{d}m(z)$ $=$ $\int_{\delta\mathbb{D}}\!\bigl{|}h^{\prime}\circ f\bigr{|}^{p}\mathbf{B}_{p}(\textup{D}f)\,\mathrm{d}m(z)$
$\geq$ $\sup_{\delta\mathbb{D}}\bigl{|}h^{\prime}\circ f\bigr{|}^{p}\int_{\delta\mathbb{D}}\!\mathbf{B}_{p}(\textup{D}f)\,\mathrm{d}m(z)$
$\geq$ $c_{p}^{p}\frac{m^{p}}{(1-\delta)^{pK}}\int_{\mathbb{D}}\!\mathbf{B}_{p}(\textup{D}f)\,\mathrm{d}m(z)$
$\geq$ $-c_{p}^{p}\frac{m^{p}}{(1-\delta)^{pK}}\pi,$
as required. ∎
3.4. Proof of Theorem 1.5
Proof.
We split the proof into two steps (building on our prerequisites) and remark that the assumption $u\in\mathrm{W}^{1,p}_{\operatorname{loc}}(\Omega)$ only is used in the second step.
By Proposition 3.6, the sequence $\bigl{(}\mathbf{B}_{p}(\textup{D}u_{j})\bigr{)}$ is bounded in $\mathrm{L}^{1}_{\operatorname{loc}}(\Omega)$. Because $\mathrm{L}^{1}_{\operatorname{loc}}(\Omega)\subset\mathrm{C}_{c}(\Omega)^{\ast}$ continuously, the Banach-Alaoglu compactness theorem implies that any subsequence of $\bigl{(}\mathbf{B}_{p}(\textup{D}u_{j})\bigr{)}$ admits a further subsequence that converges $\mbox{}^{\ast}$ to a Radon measure on $\Omega$. Any such limit must necessarily be a nonpositive Radon measure. Assume that $-\lambda$ is such a $\mbox{}^{\ast}$ limit and consider the Lebesgue-Radon-Nikodym decomposition
$\lambda=\frac{\mathrm{d}\lambda}{\mathrm{d}\mathscr{L}^{2}}\mathscr{L}^{2}+\lambda^{s},\,\mbox{ where }\lambda^{s}\perp\mathscr{L}^{2}.$
We next estimate each term in this decomposition.
Step 1. $-\frac{\mathrm{d}\lambda}{\mathrm{d}\mathscr{L}^{2}}\geq\mathbf{B}_{p}(\textup{D}u)$ holds $\mathscr{L}^{2}$ almost everywhere in $\Omega$. The proof consist of a rather general localization argument at a point of differentiabilty of $u$. Arguments of this sort are known in the literature as blow-up arguments and in the present context go back to I. Fonseca and S. Müller [25]. The fact that the Burkholder functional is nonpositive on its effective domain means that one must be more careful than in the classical set-up, where the considered functionals are bounded below and therefore concentration effects are not so harmful.
Let $z_{0}\in\Omega$ be a point satisfying
(3.16) $u\mbox{ is differentiable at }z_{0}\mbox{ with Jacobi matrix }A_{0}\equiv\textup{D}u(z_{0})$
and
(3.17) $\frac{\mathrm{d}\lambda}{\mathrm{d}\mathscr{L}^{2}}(z_{0})=\lim_{r\searrow 0}\frac{\lambda(\mathbb{D}_{r}(z_{0}))}{\mathscr{L}^{2}(\mathbb{D}_{r}(z_{0}))}.$
From the local uniform convergence of $u_{j}$ to $u$ and (3.16) follows that
$\lim_{r\searrow 0}\lim_{j\to\infty}\frac{1}{r}\sup_{z\in\mathbb{D}_{r}(z_{0})}\bigl{|}u_{j}(z)-u_{j}(z_{0})-A_{0}(z-z_{0})\bigr{|}=0.$
Next, for $r\in(0,\operatorname{dist}(z_{0},\partial\Omega))$ with $\lambda(\partial\mathbb{D}_{r}(z_{0}))=0$ we have
$\lim_{j\to\infty}\int_{\mathbb{D}_{r}(z_{0})}\!\mathbf{B}_{p}(\textup{D}u_{j})\,\mathrm{d}m(z)=-\lambda(\mathbb{D}_{r}(z_{0})),$
and so by (3.17) we get in particular
$\lim_{E\notowner r\searrow 0}\lim_{j\to\infty}\mathchoice{{\vbox{\hbox{$\textstyle-$}}\kern-4.86108pt}}{{\vbox{\hbox{$\scriptstyle-$}}\kern-3.25pt}}{{\vbox{\hbox{$\scriptscriptstyle-$}}\kern-2.29166pt}}{{\vbox{\hbox{$\scriptscriptstyle-$}}\kern-1.875pt}}\!\int_{\mathbb{D}_{r}(z_{0})}\!\mathbf{B}_{p}(\textup{D}u_{j})\,\mathrm{d}m(z)=-\frac{\mathrm{d}\lambda}{\mathrm{d}\mathscr{L}^{2}}(z_{0}),$
where $E\equiv\{r\in(0,\operatorname{dist}(z_{0},\partial\Omega)):\,\lambda(\partial\mathbb{D}_{r}(z_{0}))>0\}$ is an at most countable set. Fix $s\in(0,1)$. In view of the above we may choose a null sequence $r_{j}\searrow 0$ such that in terms of
$v_{j}(z)\equiv\frac{u_{j}(z_{0}+r_{j}z)-u_{j}(z_{0})}{r_{j}},\,z\in\mathbb{D},$
we have
$\lim_{j\to\infty}\int_{\mathbb{D}_{s}(0)}\!\mathbf{B}_{p}(\textup{D}v_{j})\,\mathrm{d}m(z)=-\pi s^{2}\frac{\mathrm{d}\lambda}{\mathrm{d}\mathscr{L}^{2}}(z_{0}).$
Here it is clear that $v_{j}$ are $K$-quasiregular and $v_{j}\to A_{0}$ uniformly on $\mathbb{D}$. The matrix $A_{0}$ is $K$-quasiconformal, so if its conformal-anticonformal coordinates are $(A_{0}^{+},A_{0}^{-})$, then $|A_{0}^{-}|\leq|A_{0}^{+}|/(p-1)$. Its Stoilow factorization is therefore (when $A_{0}\neq 0$) $A_{0}=h\circ f$ with $h(z)\equiv A_{0}^{+}z$ and
$f(z)\equiv\left\{\begin{array}[]{cl}z+\frac{A_{0}^{-}}{A_{0}^{+}}\bar{z}&\mbox{ if }|z|\leq 1,\\ z+\frac{A_{0}^{-}}{A_{0}^{+}}\frac{1}{z}&\mbox{ if }|z|>1.\end{array}\right.$
If $v_{j}=h_{j}\circ f_{j}$ is the Stoilow factorization of $v_{j}$, then using uniqueness of the factorization it is not difficult to see that $f_{j}\to f$ uniformly on $\mathbb{C}$ (if $A_{0}=0$ we use normality to extract a convergent subsequence and this then defines $f$) and $h_{j}\to h$ locally uniformly on $f(\mathbb{D})$. Consequently we have that $h_{j}^{\prime}(f_{j}(z))\to A_{0}^{+}$ uniformly in $z\in\mathbb{D}_{s}(0)$, hence using the properties of $\mathbf{B}_{p}$ and in particular the Burkholder area inequality (see [4]) and that $A_{F_{j}}\to(1,\frac{A_{0}^{-}}{A_{0}^{+}})$:
$-\pi s^{2}\frac{\mathrm{d}\lambda}{\mathrm{d}\mathscr{L}^{2}}(z_{0})$ $=$ $\lim_{j\to\infty}\int_{\mathbb{D}_{s}(0)}\!\mathbf{B}_{p}(\textup{D}v_{j})\,\mathrm{d}m(z)$
$=$ $\lim_{j\to\infty}\int_{\mathbb{D}_{s}(0)}\!\bigl{|}h_{j}^{\prime}\circ f_{j}\bigr{|}^{p}\mathbf{B}_{p}(\textup{D}f_{j})\,\mathrm{d}m(z)$
$=$ $\lim_{j\to\infty}\int_{\mathbb{D}_{s}(0)}\!|A_{0}^{+}|^{p}\mathbf{B}_{p}(\textup{D}f_{j})\,\mathrm{d}m(z)$
$\geq$ $|A_{0}^{+}|^{p}\limsup_{j\to\infty}\int_{\mathbb{D}}\!\mathbf{B}_{p}(\textup{D}f_{j})\,\mathrm{d}m(z)$
$\geq$ $|A_{0}^{+}|^{p}\limsup_{j\to\infty}\bigl{(}\pi\mathbf{B}_{p}(A_{f_{j}})\bigr{)}=\pi\mathbf{B}_{p}(A_{0}).$
Finally, taking $s\nearrow 1$ we conclude that $-\frac{\mathrm{d}\lambda}{\mathrm{d}\mathscr{L}^{2}}(z_{0})\geq\mathbf{B}_{p}(\textup{D}u(z_{0}))$ holds at all points $z_{0}\in\Omega$ satisfying (3.16) and (3.17). Because this includes $\mathscr{L}^{2}$ almost all points in $\Omega$ the proof of Step 1 is finished.
Step 2. $\lambda^{s}=0$ when $u\in\mathrm{W}^{1,p}_{\operatorname{loc}}(\Omega)$.
It suffices to show that $\lambda^{s}(B)=0$ for each disk $B=\mathbb{D}_{R}(w_{0})\Subset\Omega$. Fix such a disk $B$. Let $z_{0}\in B$ satisfy
(3.18) $\lim_{r\searrow 0}\frac{\lambda(\mathbb{D}_{r}(z_{0}))}{\mathscr{L}^{2}(\mathbb{D}_{r}(z_{0}))}=+\infty$
and
(3.19) $\lim_{r\searrow 0}\frac{1}{\lambda(\mathbb{D}_{r}(z_{0}))}\int_{\mathbb{D}_{r}(z_{0})}\!|\textup{D}u|^{p}\,\mathrm{d}m(z)=0.$
Put
$v_{j,r}(z)\equiv\frac{u_{j}(z_{0}+rz)-u_{j}(z_{0})}{r\rho_{r}},\quad z\in\mathbb{D},\mbox{ where }\rho_{r}\equiv\left(\frac{\lambda(\mathbb{D}_{r}(z_{0}))}{\mathscr{L}^{2}(\mathbb{D}_{r}(z_{0}))}\right)^{\frac{1}{p}}.$
We invoke (3.19) as follows. Recall that $u_{j}$ are $K$-quasiregular and that $u_{j}\to u$ locally uniformly on $\Omega$ and boundedly in $\mathrm{W}^{1,q}_{\operatorname{loc}}(\Omega)$ for each $q<p$, hence we get as $j\to\infty$:
$\int_{\mathbb{D}}\!|\textup{D}v_{j,r}|^{2}\,\mathrm{d}m(z)$ $=$ $\frac{\pi}{\rho_{r}^{2}}\mathchoice{{\vbox{\hbox{$\textstyle-$}}\kern-4.86108pt}}{{\vbox{\hbox{$\scriptstyle-$}}\kern-3.25pt}}{{\vbox{\hbox{$\scriptscriptstyle-$}}\kern-2.29166pt}}{{\vbox{\hbox{$\scriptscriptstyle-$}}\kern-1.875pt}}\!\int_{\mathbb{D}_{r}(z_{0})}\!|\textup{D}u_{j}|^{2}\,\mathrm{d}m(z)\leq\frac{\pi K}{\rho_{r}^{2}}\mathchoice{{\vbox{\hbox{$\textstyle-$}}\kern-4.86108pt}}{{\vbox{\hbox{$\scriptstyle-$}}\kern-3.25pt}}{{\vbox{\hbox{$\scriptscriptstyle-$}}\kern-2.29166pt}}{{\vbox{\hbox{$\scriptscriptstyle-$}}\kern-1.875pt}}\!\int_{\mathbb{D}_{r}(z_{0})}\!\det\textup{D}u_{j}\,\mathrm{d}m(z)$
$\to$ $\frac{\pi K}{\rho_{r}^{2}}\mathchoice{{\vbox{\hbox{$\textstyle-$}}\kern-4.86108pt}}{{\vbox{\hbox{$\scriptstyle-$}}\kern-3.25pt}}{{\vbox{\hbox{$\scriptscriptstyle-$}}\kern-2.29166pt}}{{\vbox{\hbox{$\scriptscriptstyle-$}}\kern-1.875pt}}\!\int_{\mathbb{D}_{r}(z_{0})}\!\det\textup{D}u\,\mathrm{d}m(z)\leq\frac{\pi K}{\rho_{r}^{2}}\mathchoice{{\vbox{\hbox{$\textstyle-$}}\kern-4.86108pt}}{{\vbox{\hbox{$\scriptstyle-$}}\kern-3.25pt}}{{\vbox{\hbox{$\scriptscriptstyle-$}}\kern-2.29166pt}}{{\vbox{\hbox{$\scriptscriptstyle-$}}\kern-1.875pt}}\!\int_{\mathbb{D}_{r}(z_{0})}\!|\textup{D}u|^{2}\,\mathrm{d}m(z)$
$\leq$ $\pi K\left(\rho_{r}^{-p}\mathchoice{{\vbox{\hbox{$\textstyle-$}}\kern-4.86108pt}}{{\vbox{\hbox{$\scriptstyle-$}}\kern-3.25pt}}{{\vbox{\hbox{$\scriptscriptstyle-$}}\kern-2.29166pt}}{{\vbox{\hbox{$\scriptscriptstyle-$}}\kern-1.875pt}}\!\int_{\mathbb{D}_{r}(z_{0})}\!|\textup{D}u|^{p}\,\mathrm{d}m(z)\right)^{\frac{2}{p}}$
$=$ $\pi K\left(\frac{1}{\lambda\bigl{(}\mathbb{D}_{r}(z_{0})\bigr{)}}\int_{\mathbb{D}_{r}(z_{0})}\!|\textup{D}u|^{p}\,\mathrm{d}m(z)\right)^{\frac{2}{p}}.$
Consequently,
(3.20) $\lim_{r\searrow 0}\lim_{j\to\infty}\int_{\mathbb{D}}\!|\textup{D}v_{j,r}|^{2}\,\mathrm{d}m(z)=0.$
Next, we get as in Step 2,
$\lim_{j\to\infty}\int_{\mathbb{D}_{r}(z_{0})}\!\mathbf{B}_{p}(\textup{D}u_{j})\,\mathrm{d}m(z)=-\lambda(\mathbb{D}_{r}(z_{0}))$
when $\lambda(\partial\mathbb{D}_{r}(z_{0}))=0$, hence
$\lim_{E\notowner r\searrow 0}\lim_{j\to\infty}\int_{\mathbb{D}}\!\mathbf{B}_{p}(\textup{D}v_{j,r})\,\mathrm{d}m(z)=-1,$
where $E\equiv\{r\in(0,\operatorname{dist}(z_{0},\partial\Omega)):\,\lambda(\partial\mathbb{D}_{r}(z_{0}))>0\}$ is an at most countable set. In order to combine this with (3.18) we require some routine arguments from measure theory that we briefly recall here for convenience of the reader. Define
$\Lambda(t)\equiv\limsup_{E\notowner r\searrow 0}\frac{\lambda\bigl{(}\mathbb{D}_{tr}(z_{0})\bigr{)}}{\lambda\bigl{(}\mathbb{D}_{r}(z_{0})\bigr{)}},\,t\in(0,1].$
We assert that (3.18) implies that $\Lambda(t)\geq t^{2}$ for all $t\in(0,1]$. Indeed suppose for a contradiction that it fails at some $s\in(0,1)$, so that $\Lambda(s)<s^{2}$. First note that since the function $r\mapsto\lambda\bigl{(}\mathbb{D}_{r}(z_{0})\bigr{)}$ is left-continuous and the set $E$ is at most countable we actually have that
$\Lambda(t)=\limsup_{r\searrow 0}\frac{\lambda\bigl{(}\mathbb{D}_{tr}(z_{0})\bigr{)}}{\lambda\bigl{(}\mathbb{D}_{r}(z_{0})\bigr{)}},\,t\in(0,1],$
and so choosing $t\in(0,s)$ with $\Lambda(s)<t^{2}$ we get by iteration that for some $\delta>0$,
$\lambda\bigl{(}\mathbb{D}_{s^{j}\delta}(z_{0})\bigr{)}<t^{2j}\lambda\bigl{(}\mathbb{D}_{\delta}(z_{0})\bigr{)}$
holds for all $j\in\mathbb{N}$. But this contradicts (3.18) since then, as $j\to\infty$,
$\frac{\lambda\bigl{(}\mathbb{D}_{s^{j}\delta}(z_{0})\bigr{)}}{s^{2j}\delta^{2}}<\left(\frac{t}{s}\right)^{2j}\frac{\lambda\bigl{(}\mathbb{D}_{\delta}(z_{0})\bigr{)}}{\delta^{2}}\to 0.$
In terms of the maps $v_{j,r}$ this amounts to that the bound
(3.21) $\limsup_{E\notowner r\searrow 0}\lim_{j\to\infty}\int_{\mathbb{D}_{s}(0)}\!\mathbf{B}_{p}(\textup{D}v_{j,r})\,\mathrm{d}m(z)\leq-\pi s^{2}$
holds for each $s\in(0,1]$. In order to conclude the proof we fix $s\in(0,1)$, use (3.20) and (3.21) to select a null sequence $r_{j}\searrow 0$ such that $v_{j}\equiv v_{j,r_{j}}\to 0$ strongly in $\mathrm{W}^{1,2}(\mathbb{D})$ and
$\lim_{j\to\infty}\int_{\mathbb{D}_{s}(0)}\!\mathbf{B}_{p}(\textup{D}v_{j})\,\mathrm{d}m(z)\leq-\pi s^{2}.$
Now $v_{j}\colon\mathbb{D}\to\mathbb{C}$ are $K$-quasiregular maps and so we may Stoilow factorize as $v_{j}=h_{j}\circ f_{j}$, where $f_{j}\colon\mathbb{C}\to\mathbb{C}$ are $K$-quasiconformal principal and $h_{j}\colon f_{j}(\mathbb{D})\to\mathbb{C}$ are holomorphic. Taking a further subsequence if necessary (not relabelled) we may assume that $f_{j}\to f$ uniformly on $\mathbb{C}$, where $f\colon\mathbb{C}\to\mathbb{C}$ is a $K$-quasiconformal principal map. We then also have that $h_{j}\to 0$ locally uniformly on $f(\mathbb{D})$. It follows that $h_{j}^{\prime}\circ f_{j}\to 0$ uniformly on $\mathbb{D}_{s}(0)$ and therefore
$-\pi s^{2}$$\geq$ $\lim_{j\to\infty}\int_{\mathbb{D}_{s}(0)}\!\bigl{|}h_{j}^{\prime}\circ f_{j}\bigr{|}^{p}\mathbf{B}_{p}(\textup{D}f_{j})\,\mathrm{d}m(z)$
$\geq$ $\sup_{\mathbb{D}_{s}(0)}\bigl{|}h_{j}^{\prime}\circ f_{j}\bigr{|}^{p}\int_{\mathbb{D}}\!\mathbf{B}_{p}(\textup{D}f_{j})\,\mathrm{d}m(z)$
$\geq$ $-\pi\sup_{\mathbb{D}_{s}(0)}\bigl{|}h_{j}^{\prime}\circ f_{j}\bigr{|}^{p}\to 0$
as $j\to\infty$, which is impossible. The set of points $z_{0}\in B$ satisfying (3.18) and (3.19) is therefore empty, and since $\lambda^{s}$ almost all points in $B$ have those properties we infer that $\lambda^{s}(B)=0$ as required. ∎
For the Dirichlet classes we have the following immediate corollary.
Corollary 3.7.
Let $\Omega$ be a bounded open subset of $\mathbb{C}$ and $g\colon\mathbb{C}\to\mathbb{C}$ be a $K$-quasiregular map. If $u_{j}\in g+\mathrm{W}^{1,p_{K}}_{0}(\Omega)$ are $K$-quasiregular and $u_{j}\to u$ weakly in $\mathrm{W}^{1,p_{K}}(\Omega)$, then $u\in g+\mathrm{W}^{1,p_{K}}_{0}(\Omega)$ is $K$-quasiregular and
$\liminf_{j\to\infty}\int_{\Omega}\!\mathbf{B}_{p_{K}}(\textup{D}u_{j})\,\mathrm{d}m(z)\geq\int_{\Omega}\!\mathbf{B}_{p_{K}}(\textup{D}u)\,\mathrm{d}m(z).$